# Fast and efficient formation of stable tetraatomic molecules from ultracold atoms via generalized stimulated Raman exact passage

### v1

@@ III.1 Model and Method
First, we postulate that $\Delta$ is the largest evolution frequency Mackie and Phou (2010); Liu et al. (2019b), $\Delta\!\gg\!\Omega_{i}(i\!=\!1,2,3,4)$. In this limit, adiabatic elimination (AE) enables the derivation of an effective three-state model, which takes the form of
@@ III.2 Result and Discussion
Figure. 6(a) illustrates the Rabi frequencies required to achieve the desired population transfer result. From this figure, it can be observed that the first and fourth Rabi frequencies not only exhibit the same temporal profile and equal magnitudes, but also straddle the middle two ones. They feature near-Gaussian profiles with sufficient smoothness, thus posing no actual implementation challenges Du et al. (2016); Zheng et al. (2022). The time evolution results from the initial state $|1\rangle$ to the target state $|5\rangle$ is displayed in Fig. 6(b). It can be observed that a complete state transfer can be achieved by employing four modified fields. A high transfer efficiency is helpful to minimize the loss in phase-space density when an ensemble of weakly-bound molecules is transferred to the ground state. In addition, we note that, throughout the entire transfer process, the intermediate state $|3\rangle$ receives some amount of transient population, while the populations of the two excited states are negligible. This process does not rely on the existence of dark states; instead, we utilize the AE protocol to achieve the stable decoupling of two excited states from dynamics. Thus, although these two states are prone to decaying out of the system, the transfer process remains scarcely swayed by them.
@@
Finally, we are going to perform the targeted numerical simulations to verify the necessity of the condition (18). Figure. 8(a) shows a contour plot of transfer efficiency against the Rabi frequency deviations $\delta\tilde{\Omega}_{1}$ and $\delta\tilde{\Omega}_{4}$ when $\tilde{\Omega}_{2}$ and $\tilde{\Omega}_{3}$ are fixed at their ideal values. Clearly, $\delta\!=\!0$ implies that an ideal operating condition is achieved; $\delta\!\neq\!0$ indicates that the condition (18) is violated. According to the results shown in this figure, any deviation from the ideal condition results in a significant drop in efficiency. In particular, when the deviation is relatively large, the transfer efficiency will suffer a loss of around $10\%$. In addition, we observe that the deviations of $\tilde{\Omega}_{1}$ and $\tilde{\Omega}_{4}$ exert symmetric effects on the transfer efficiency. This confirms that this condition is an indispensable prerequisite for obtaining high population transfer efficiency. The transfer efficiency can remain at a high level if and only if $\tilde{\Omega}_{1}\!=\!\tilde{\Omega}_{4}$ and they should exactly match the ideal value.
Figure. 8(b) presents a contour plot of the transfer efficiency as a function of $\delta\tilde{\Omega}_{2}$ and $\delta\tilde{\Omega}_{3}$ when condition (18) is fully satisfied. In this case, $\tilde{\Omega}_{1}$ and $\tilde{\Omega}_{4}$ will passively vary with the other two ones. It is clear that the transfer efficiency can keep at a high level, even when the deviation of the Rabi frequency deviations within $|\delta\tilde{\Omega}_{2}/\tilde{\Omega}_{2}|\!=\!|\delta\tilde{\Omega}_{3}/\tilde{\Omega}_{3}|\!=\!10\%$. This result further demonstrates the necessity of the condition (18).
@@ IV Conclusions and Outlooks
To conclude, we present a theoretical formalism for the formation of stable ultracold tetraatomic molecules. The process begins with a mixture of ultracold atoms: first, generalized nonlinear STIREP is employed to coherently assemble atoms into tetraatomic molecules; subsequently, to ensure the resulting molecules exhibit sufficient stability, we propose that the C-STIREP technique be used to transfer them to the stable ground state. Numerical analyses show that the proposed two-step strategy enables the production of stable ultracold molecules with nearly ideal efficiency, while featuring excellent accelerability and high robustness.
The main advantages of our work are organized as follows: (i) It paves a potential pathway for forming stable ultracold tetraatomic molecules, and in principle, has the potential to be extended to preparing larger polyatomic molecular complexes Dou et al. (2013); (ii) Unlike traditional methods, the adiabatic condition is released in our approach, thus enabling fast speed and high efficiency. Additionally, the transient populations of short-lived states can be effectively suppressed without relying on the dark state; (iii) Although the protocol follows the spirit of STA, it does not introduce extra couplings, and the applied pulses pose no actual implementation difficulties; (iv) Numerous exact solutions support the achievement of predefined goals, and the above-discussed examples (7) and (20) are not exhaustive Chen and Muga (2012); Kang et al. (2016); Baksic et al. (2016); Wu et al. (2019).
Finally, we point out that the tetraatomic molecule preparation process suggested in this study aligns closely with the standard procedure for diatomic molecule, with both following the similar route of “first creating unstable molecular samples by assembling pre-cooled atom one by one, and then converting them to sufficiently stable ground state”. Inspired by the studies in Refs. Mackie and Phou (2010); Price and Yelin (2019), we will, in the future, commit to exploring a coherent strategy that starts directly from ultracold atoms and enables the robust, fast and efficient production of stable UPMs. This problem may be worth mentioning.

### v3

@@ III.1 Model and Method
First, we postulate that $\Delta$ is the largest evolution frequency Lugiato et al. (1984); Mackie and Phou (2010); Liu et al. (2019b), $\Delta\!\gg\!\Omega_{i}(i\!=\!1,2,3,4)$. In this limit, adiabatic elimination (AE) enables the derivation of an effective three-state model, which takes the form of
@@ III.2 Result and Discussion
Figure. 6(a) illustrates the Rabi frequencies required to achieve the desired population transfer result. From this figure, it can be observed that the first and fourth Rabi frequencies not only exhibit the same temporal profile and equal magnitudes, but also straddle the middle two ones. They feature near-Gaussian profiles with sufficient smoothness, thus posing no actual implementation challenges Du et al. (2016); Zheng et al. (2022); Norambuena et al. (2024). The time evolution results from the initial state $|1\rangle$ to the target state $|5\rangle$ is displayed in Fig. 6(b). It can be observed that a complete state transfer can be achieved by employing four modified fields. A high transfer efficiency is helpful to minimize the loss in phase-space density when an ensemble of weakly-bound molecules is transferred to the ground state. In addition, we note that, throughout the entire transfer process, the intermediate state $|3\rangle$ receives some amount of transient population, while the populations of the two excited states are negligible. This process does not rely on the existence of dark states; instead, we utilize the AE protocol to achieve the stable decoupling of two excited states from dynamics. Thus, although these two states are prone to decaying out of the system, the transfer process remains scarcely swayed by them.
@@
Finally, we are going to perform the targeted numerical simulations to verify the necessity of Condition (18). Figure. 8(a) shows a contour plot of transfer efficiency against the Rabi frequency deviations $\delta\tilde{\Omega}_{1}$ and $\delta\tilde{\Omega}_{4}$ when $\tilde{\Omega}_{2}$ and $\tilde{\Omega}_{3}$ are fixed at their ideal values. Clearly, $\delta\!=\!0$ implies that an ideal operating condition is achieved; $\delta\!\neq\!0$ indicates that the Condition (18) is violated. According to the results shown in this figure, we can find that any deviation from the ideal condition will result in a significant drop in efficiency. In particular, when the deviation is relatively large, the transfer efficiency will suffer a loss of around $10\%$. Except for this, we can observe that the deviations $\tilde{\Omega}_{1}$ and $\tilde{\Omega}_{4}$ exert symmetric effects on the transfer efficiency. This confirms that this condition is an indispensable prerequisite for obtaining high population transfer efficiency. The transfer efficiency can remain at a high level if and only if $\tilde{\Omega}_{1}\!=\!\tilde{\Omega}_{4}$ and they should exactly match the ideal value.
On the other hand, Figure. 8(b) presents a contour plot of the transfer efficiency as a function of $\delta\tilde{\Omega}_{2}$ and $\delta\tilde{\Omega}_{3}$, with Condition (18) maintained; in this case, $\tilde{\Omega}_{1}$ and $\tilde{\Omega}_{4}$ will passively vary with the other two ones. It is clear that the transfer efficiency can keep at a high level, even when the deviation of the Rabi frequency deviations within $|\delta\tilde{\Omega}_{2}/\tilde{\Omega}_{2}|\!=\!|\delta\tilde{\Omega}_{3}/\tilde{\Omega}_{3}|\!=\!10\%$. This result further demonstrates the necessity of Condition (18).
@@ IV Conclusions and Outlooks
To conclude, we present a theoretical formalism for the formation of stable ultracold tetratomic molecules. The process begins with a mixture of ultracold atoms: first, generalized nonlinear STIREP is employed to coherently assemble atoms into tetratomic molecules; subsequently, to ensure the resulting molecules exhibit sufficient stability, we propose that the C-STIREP technique be used to transfer them to the stable ground state. Numerical analyses show that the proposed two-step strategy enables the production of stable ultracold molecules with nearly ideal efficiency, while featuring excellent accelerability and high robustness.
The main advantages of our work are organized as follows: (i) It paves a potential pathway for forming stable ultracold tetratomic molecules, and in principle, has the potential to be extended to preparing larger polyatomic molecular complexes Dou et al. (2013); (ii) Unlike traditional methods, the adiabatic condition is released in our approach, thus enabling fast speed and high efficiency. Additionally, the transient populations of short-lived states can be effectively suppressed without relying on the dark state; (iii) Although the protocol follows the spirit of STA, it does not introduce extra couplings, and the applied pulses pose no actual implementation difficulties; (iv) Numerous exact solutions support the achievement of predefined goals, and the above-discussed examples (7) and (20) are not exhaustive Chen and Muga (2012); Kang et al. (2016); Baksic et al. (2016); Wu et al. (2019).
Finally, we point out that the tetratomic molecule preparation process suggested in this study aligns closely with the standard procedure for diatomic molecule, with both following the similar route of “first creating unstable molecular samples by assembling pre-cooled atom, and then converting them to sufficiently stable ground state”. Inspired by the studies in Refs. Mackie and Phou (2010); Price and Yelin (2019), we will, in the future, commit to exploring a coherent strategy that starts directly from ultracold atoms and enables the robust, fast and efficient production of stable UPMs. This problem may be worth mentioning.